OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = 5/2 x + 2.

OpenStudy (anonymous):

@Kitt020912 @AriPotta

OpenStudy (anonymous):

Ooh perpendicular equation.. I got this(:

OpenStudy (aripotta):

lines that are perpendicular have negative reciprocal slopes

OpenStudy (aripotta):

so the new line would have a slope of -2/5

OpenStudy (anonymous):

Oh yes, the explaining would help me majorly as well - I have several more of these types of questions :3 ♥

OpenStudy (anonymous):

right :D

OpenStudy (aripotta):

and we can put this information into point-slope form

OpenStudy (anonymous):

& I would like to be able to understand them, lol

OpenStudy (aripotta):

y - y1 = m(x - x1) so we're given this point, (-5, -1) and we can plop that in the equation along with the slope y + 1 = -2/5(x + 5) now put that into slope-intercept form, sam

OpenStudy (aripotta):

slope-intercept form being y = mx + b first you gotta distribute that -2/5

OpenStudy (anonymous):

ari wins, I'm no good at explaining this >_<

OpenStudy (anonymous):

Ok, so you would have: y + 1 = -2/5m - 2 That's what as far as distributing the -2/5 and I have a feeling thats wrong...

OpenStudy (aripotta):